OpenStudy (anonymous):

If the diameterof the plunger is 5 cm, what is the area of the plunger?

OpenStudy (anonymous):

math question. you'd get an answer faster if you asked it in the right section.

OpenStudy (anonymous):

A= pi * r^2 = 3,14 * (2,5cm)^2 = 19,625 cm^2